Two missiles speed directly towards each other, one at 12000 km per hour and the other at 18000 km
per hour. How far apart they would be exactly 30 seconds before they collide?

Answer by stanbon(75887)

Two missiles speed directly towards each other, one at 12000 km per hour and the other at 18000 km
per hour. How far apart would they be exactly 30 seconds before they collide?
-------
Note:: 30 seconds = 30/3600 = 1/120 hour
---------
Closing speed of the missiles is 12000+18000 = 30,000 kph
------
Distance covered in 30 seconds = (1/12)(30,000) = 250 km
=========================
Cheers,
Stan H.
=========================